OpenStudy (anonymous):

Jack wants to find the solution to a system of equations using y = x + 3 and y = x2 + 6x - 21. Betty says that Jack can solve x2 + 5x - 24 = 0 to find the x-coordinates of the solutions of the system. Explain and demonstrate why Betty is correct.

OpenStudy (anonymous):

The intersection of the line \(x+3\) and the parabola \(x^2+6x-21\) occurs when the x and y coordinates are the same, which means you can set the y's equal to each other: \[x+3=x^2+6x-21\] Rearranging, you set equal to zero: \[x^2+5x-24=0\]